Assume that the salaries of elementary school teachers in the unaired states have a mean of $32,000 and a standard deviation of $3000. if 100 teachers are randomly selected, would it be unusual to obtain a sample mean of 32450? answer by calculating the appropriate z-score.
Answer #1

Solution :

Given that ,

mean = \mu = 32000

standard deviation = \sigma = 3000

n = 100

\mu\bar x = 32000

\sigma\bar x =\sigma  / \sqrt n = 3000 / \sqrt 100=300

\bar x =32450

using z score formula

z=(\bar x - \mu \bar x ) / \sigma \bar x

= 32450-32000/300

=1.5

z score =1.5